OpenStudy (anonymous):

how to find the volume of the solid generated by rotating the region bounded by y=x, y=x^1/2) about y=1

OpenStudy (anonymous):

*y=x^(1/2)

OpenStudy (anonymous):

assuming y = x^1/2 to be the function. I don't think you can find the volume using standard methods because the y starts from 0 and there would be overlaps. If the axis of revolution was y=0, or if you had something like y=x^(1/2)+1, it would be possible.

OpenStudy (anonymous):

thank you